Find the ratio of:
 ₹ 5 to 50 paise
Find the ratio of: 
 ₹ 5 to 50 paise

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that,
₹1 = 100 paise Then,
₹5 = 5 × 100 = 500 paise
Now we have to find the ratio,
= 500/50
= 10/1
So, the required ratio is 10: 1.

Thanks